OpenStudy (anonymous):

Which unit would be most appropriate for measuring the width of a house? A. km B. mm C. m D. cm

OpenStudy (zaberangelr):

a or c

OpenStudy (zaberangelr):

gada think how long is a kilometer

OpenStudy (zaberangelr):

and how long is a meter

OpenStudy (alexandervonhumboldt2):

km is 1000 meters it is too long

OpenStudy (zaberangelr):

so its c

OpenStudy (alexandervonhumboldt2):

km are used to measure road distance

OpenStudy (zaberangelr):

boom

OpenStudy (zaberangelr):

cm is small

OpenStudy (zaberangelr):

mm is smaler then cm

OpenStudy (anonymous):

thanks

OpenStudy (zaberangelr):

np

OpenStudy (zaberangelr):

hello k12

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Unit 2 | Lesson 15 Unit Checkpoint: Distance: Addition Equations Question Navigator Choose the answer. 12. The perimeter of this rectangle can be calculated using the equation P = 2 × 18.5 + 2 × 7.2. Which other equation could be used to calculate the perimeter? A. P = 2 (18.5 + 7.2) B. P = 2 × 18.5 × 7.2 C. P = (2 • 18.5) + 7.2 D. P = 2 × 18.5 + 7.2 Question Resources

OpenStudy (anonymous):

i will fan you

OpenStudy (anonymous):

A rectangle has a length of 10 inches and a perimeter of 48 inches. Which equation could be used to find the width of the rectangle? A. 48 = 10w B. 48 = 20 – 2w C. 48 = 40w D. 48 = 20 + 2w

OpenStudy (anonymous):

9. (3 pt) What is the perimeter of this rectangular field? m Rectangle 93.2 m by 78.6 m

OpenStudy (anonymous):

help help help help help...

OpenStudy (anonymous):

Jason's dad set the trip meter in their car to 0 miles as they left Jacksonville, Florida, on their way to Houston, Texas. When they stopped in New Orleans, Louisiana, Jason saw that the trip meter read 507.8 miles. When they arrived in Houston it read 824.5 miles. What is the distance from New Orleans to Houston? A. 326.7 miles B. 316.7 miles C. 1332.3 miles D. 1333.2 miles

OpenStudy (anonymous):

i will friend you and medal

OpenStudy (anonymous):

Julie ran 32.4 miles in May and 35.8 miles in June. Greg ran the same total distance as Julie, but only ran 29.6 miles in May. How far did Greg run in June? A. 33.0 miles B. 38.6 miles C. 2.8 miles D. 6.2 miles

OpenStudy (anonymous):

A rectangle has a length of 10 inches and a perimeter of 48 inches. Which equation could be used to find the width of the rectangle? A. 48 = 20 + 2w B. 48 = 20 – 2w C. 48 = 40w D. 48 = 10w

OpenStudy (anonymous):

Anne walked 3.2 kilometers. How many meters did Anne walk? A. 3200 meters B. 320 meters C. 0.032 meters D. 0.0032 meters